Golden Ring Company produces two types of product: Large and Larger. Two work orders for two batches of the products are shown below, along with some additional cost information:

Large Larger
Work Order 10 Work Order 11
Direct materials (actual costs) $45,000 $75,000
Applied conversion costs:
Mixing ? ?
Cooking $12,000 $12,000
Bottling $10,000 $15,000
Batch size (bottles) 5,000 5,000


In the Mixing Department, conversion costs are applied on the basis of direct labor hours. Budgeted conversion costs for the department for the year were $50,000 for labor and $125,000 for overhead. Budgeted direct labor hours were 2,500. It takes three minutes to mix the ingredients needed for each bottle.

Large (Work Order 10) and Larger (Work Order 11) flow through the Mixing Department first, then through the Cooking and Bottling departments.

What is Golden Ring Company's unit cost of Large?

a.$23.90

b.$3.50

c.$70.00

d.$16.90
Answer #1

Conversion cost rate = (50000+125000)/2500 = 70 per hour

Large mixing department cost = 70*5000*3/60 = 17500

Unit cost of large = (45000+12000+10000+17500)/5000 = 16.90

so answer is d) $16.90